\begin{thebibliography}{99}

\bibitem{Calmetetal04}  Calmet,~X., Graesser,~M. and Hsu,~S.~D.~H.,  
``Minimum length from quantum mechanics and classical general relativity,'' {\em Phys. Rev. Lett.}\/ {\bf 93}, 211101 (2004);arXiv:hep-th/0405033.%}%end \textcolor{teal}{

\bibitem{Hossenfelder13} S.~Hossenfelder,
``Minimal length scale scenarios for quantum gravity,''
{\em Living Rev. Relativity},\/ {\bf 16}, 2 (2013).

\bibitem{BHZ05} R.~ V.~Buniy, S.~D.~H.~Hsu and  A.~Zee, ``Is Hilbert space discrete?'' 
{\em Phys. Lett. B}\/ {\bf 630}, 68-72 (2005); arXiv:hep-th/0508039.


\bibitem{FaizalKhalilDas16}M.~Faizal, M.~M.~Khalil and S.~Das, ``Time crystals from minimum uncertainty,'' {\em Eur. Phys. J. C}\/ {\bf 76}, 30 \textcolor{teal}{(2016)};
arXiv:1501.03111.

\bibitem{CalmetHsu21}X.~Calmet and S.~D.~H.~Hsu, ``Fundamental limit on angular measurements and rotations from quantum mechanics and general relativity,'' arXiv:2108.11990 (2021).

\bibitem{DeWitt70}B.~S.~DeWitt, ``Quantum mechanics and reality,'' {\em Physics Today}\/ {\bf 23}, 30-35 (1970).Reprinted in \cite{DeWittGraham73}.

\bibitem{Everett57}H.~Everett~III, `` `Relative state' formulation of quantum 
mechanics,'' {\em Rev. Mod. Phys.}\/ {\bf 29}  454-462 (1957). Reprinted in \cite{DeWittGraham73}.

\bibitem{DeWittGraham73}B.~S.~DeWitt and 
N.~Graham, eds., {\em The Many Worlds Interpretation of Quantum Mechanics}
(Princeton University Press, Princeton, NJ, 1973). %end \textcolor{teal}{

\bibitem{Price95}  M.~C.~Price, ``The Everett FAQ,'' https://www.hedweb.com/everett/ (1995). 

\bibitem{Barrett99}J.~A.~Barrett, {\em The quantum mechanics of minds and worlds}\/
(Oxford University Press, Oxford,1999). 


\bibitem{Vaidman02} L.~Vaidman, ``Many-worlds interpretation of quantum mechanics,'' {\em The Stanford Encyclopedia of Philosophy}\/ (Fall 2018 Edition), E.~N.~Zalta, ed., https://plato.stanford.edu/archives/fall2018/entries/qm-manyworlds/.


\bibitem{HewittHorsman09}C.~Hewitt-Horsman, ``An introduction to many worlds in quantum computation,'' {\em Found. Phys.}\/ {\bf 39},
826-902 (2009); arXiv:0802.2504.

\bibitem{Saundersetal10}S.~Saunders, J.~Barrett, A.~Kent and D.~Wallace, eds., {\em Many Worlds? Everett, Quantum Theory \& Reality}\/, (Oxford University Press, Oxford, 2010).

\bibitem{Wallace12}D.~Wallace,  {\em The Emergent Multiverse: Quantum Theory According to the Everett Interpretation}\/, (Oxford University Press, Oxford, 2012).

\bibitem{Wallace13}D.~Wallace, ``The Everett interpretation,'' in R.~Batterman, ed., {\em The Oxford Handbook of Philosophy of Physics}\/ 
(Oxford University Press, Oxford, 2013).


\bibitem{BHZ06} R.~ V.~Buniy, S.~D.~H.~Hsu and  A.~Zee, ``Discreteness and the origin of probability in quantum mechanics,''
{\em Phys.Lett. B}\/ {\bf 640}, 219-223 (2006) ;  arXiv:hep-th/0606062.

\bibitem{Hsu12}S.~D.~H.~Hsu, ``On the origin of probability in quantum mechanics,'' {\em Mod. Phys. Lett. A}\/ {\bf 27}, 1230014 (2012); 
arXiv:1110.0549.

\bibitem{Hsu17}S.~D.~H.~Hsu, ``The measure problem in no-collapse (many worlds) quantum mechanics,'' {\em Int. J. Mod. Phys. D}\/ {\bf 26}, 1730008 (2017); arXiv:1511.08881.

\bibitem{Hsu21} S.~D.~H.~Hsu,  ``Discrete Hilbert space, the Born rule, and quantum gravity,'' {\em Mod. Phys. Lett. A}\/ {\bf 36}, 2150013 (2021); arXiv:2007.12938.

\bibitem{Rubin21} M.~A.~Rubin, ``Probability, preclusion and biological evolution in Heisenberg-picture Everett quantum mechanics,'' {\em 
Int. J. Mod. Phys. A}\/ {\bf 36}, 2150117 (2021); arXiv:2011.10029.

\bibitem{MohanPati20}B.~Mohan and A.~K.~Pati, ``Reverse quantum speed limit: How slow quantum battery can discharge?'' arXiv:2006:14523 (2020).

%\bibitem{dEspagnat76}\textcolor{teal}{B.~d'Espagnat, {\em Conceptual Foundations of Quantum Mechanics,}\/ 2nd ed.
%(W.~A.~Benjamin, Inc., Reading, MA, 1976) .}%end \textcolor{teal}{

%\textcolor{orange}{
%\bibitem{Pati95}A.~K.~Pati, ``Geometric aspects of noncyclic quantum evolutions,'' {\em Phys. Rev. A}\/ {\bf 52}, 2576-2584 (1995).
%}%end \textcolor{orange}{
%
% Everett57, DeWittGraham73, Price95, Barrett99, Vaidman02,  HewittHorsman09, Saundersetal10, Wallace12, Wallace13


% Everett57, DeWittGraham73, Price95, Barrett99, Vaidman02,  HewittHorsman09, Saundersetal10, Wallace12, Wallace13

% Everett57,DeWitt72,Graham73,Okhuwa93,Kent90,Barnumetal00,Rae09,Saunders10,Papineau10,Wallace10,GreavesMyrvold10,Kent10,Albert10,Price10,Zurek10,Schack10,Vaidman11,SebensCarroll18,

\bibitem{DeWitt72} B.~S.~DeWitt, ``The many-universes interpretation of
quantum mechanics,'' in {\em Proceedings of the International
School of Physics ``Enrico Fermi'' Course IL: Foundations of Quantum Mechanics},\/ (Academic
Press, Inc., New York, 1972). Reprinted in \cite{DeWittGraham73}

\bibitem{Graham73}N.~Graham, ``The measurement of relative frequency,'' in \cite{DeWittGraham73}.

\bibitem{Okhuwa93}Y.~Okhuwa, ``Decoherence functional and probability interpretation,''
{\em Phys. Rev. }\/ {\bf D48}, 1781-1784 (1993).

\bibitem{Kent90} A.~Kent, ``Against many-worlds interpretations,'' {\em Int. J. Mod. Phys.}\/ {\bf A5} 1745 (1990); arXiv:gr-qc/9703089.

\bibitem{Deutsch99}D.~Deutsch, ``Quantum theory of probability and decisions,'' {\em Proc.
Roy. Soc. Lond. A}\/ {\bf 455}, 3129-3137 (1999); quant-ph/9906015.

\bibitem{Barnumetal00}H.~Barnum, C.~M.~Caves, J.~Finkelstein, C.~A.~Fuchs and R.~Schack, ``Quantum probability from decision theory?'' {\em Proc. Roy. Soc. Lond.} {\bf A456}, 1175-1182 (2000); arXiv:quant-ph/9907024


\bibitem{Rae09}A.~I.~M.~Rae, ``Everett and the Born rule,'' {\em Studies in History and Philosophy of Modern Physics} {\bf 40} 243-250 
(2009); arXiv:0810.2657.

\bibitem{Saunders10} S.~Saunders, ``Chance in the Everett interpretation,'' in \cite{Saundersetal10}, 181-205.

\bibitem{Papineau10} D.~Papineau, ``A fair deal for Everettians,'' in  \cite{Saundersetal10}, 206-226.

\bibitem{Wallace10} D.~Wallace, ``How to prove the Born rule,'' in  \cite{Saundersetal10}, 227-263.

\bibitem{GreavesMyrvold10} H.~Greaves and W.~Myrvold, ``Everett and evidence,'' in  \cite{Saundersetal10}, 264-304.

\bibitem{Kent10} A.~Kent, ``One world versus many: The inadequacy of Everettian accounts of evolution, probability, and scientific confirmation,''  in  \cite{Saundersetal10}, 307-334.

\bibitem{Albert10} D.~Albert, ``Probability in the Everett interpretation,''  in  \cite{Saundersetal10}, 355-368.

\bibitem{Price10} H.~Price, ``Decisions, decisions, decisions: Can Savage salvage Everettian probability?,''  in  \cite{Saundersetal10}, 369-390.

\bibitem{Zurek10} W.~H.~Zurek, ``Quantum jumps, Born's rule, and objective reality,'' in  \cite{Saundersetal10}, 409-432.

\bibitem{Schack10} R.~Schack, ``The principal principle and probability in the many-worlds interpretation,''  in  \cite{Saundersetal10}, 467-475.

\bibitem{Vaidman11}
L.~Vaidman
``Probability and the Many-Worlds interpretation of quantum theory, ''
in {\em Proceedings of the Conference "Quantum Theory, Reconsideration of Foundations" : V\"{a}xj\"{o} (Smaland), Sweden, 17-21 June, 2001}
(V\"{a}xj\"{o} University Press, V\"{a}xj\"{o}, Sweden, 2001)); arXiv:quant-ph/0111072

\bibitem{SebensCarroll18}C.~T.~Sebens and S.~M.~Carroll,
``Self-locating uncertainty and the origin of probability in Everettian quantum mechanics.'' {\em The British Journal for the Philosophy of Science}\/, {\bf 69}, 25-74 (2018);arXiv:1405.7577.

\bibitem{Ballentine98}L.~E.~Ballentine, {\em Quantum Mechanics: A Modern Development}\/ 
(World Scientific Publishing Co. Pte. Ltd., Singapore, 1978).%end 

%
%\bibitem{Page82}\textcolor{teal}{D.~N.~Page, ``The Einstein-Podolsky-Rosen physical reality is completely described by quantum mechanics,''  {\em Phys. Lett.} {\bf  A91}, 57-60 (1982).}% end \textcolor{teal}{
%
%\bibitem{Stapp85}\textcolor{teal}{ H.~P.~Stapp, ``Bell's theorem and the foundations of quantum physics,''
%{\em Am. J. Phys.}\/ {\bf 53}, 306 (1985).}% end \textcolor{teal}{
%
%\bibitem{Tipler00} \textcolor{teal}{F.~J.~Tipler, ``Does quantum nonlocality exist? Bell's theorem and the many-worlds interpretation,'' arXiv:quant-ph/0003146 (2000).}% end \textcolor{teal}{
%
%\bibitem{Rubin01} \textcolor{teal}{ M.~A.~Rubin,  ``Locality in the Everett interpretation of 
%Heisenberg-picture quantum mechanics,'' {\em Found. Phys. Lett.}\/ {\bf 14}, 301-322 (2001);
%quant-ph/0103079.}% end \textcolor{teal}{
%
%
%\bibitem{Bacciagaluppi02} \textcolor{teal}{G.~Bacciagaluppi, ``Remarks on space-time and locality in Everett's interpretation.,'' in  
%T.~Placek and J.~Butterfield, {\em Non-locality and Modality}\/ (Kluwer Academic Publishers, Dordrecht, 2002); philsci-archive.pitt.edu/id/eprint/504. }% end \textcolor{teal}{
%
%\bibitem{TimpsonBrown02}\textcolor{teal}{ C.~G.~Timpson and  H.~R.~Brown, ``Entanglement and relativity,'' in   R.~Lupacchini and V.~Fano,  eds., {\em Understanding Physical Knowledge}\/ (CLUEB, Bologna, 2002).}% end \textcolor{teal}{

\bibitem{Rubin11}M.~A.~Rubin, ``Observers and locality in Everett quantum field theory,'' 
{\em Found.Phys.}\/ {\bf 41}, 1236-1262 (2011);arXiv:0909.2673.% end \textcolor{teal}{

%\bibitem{Vaidman15} \textcolor{teal}{L.~Vaidman, ``The Bell inequality and the many-worlds interpretation,'' in M.~Bell and
%S.~Gao, eds., {\em Quantum Nonlocality and Reality: 
%50 Years of Bell's Theorem}\/ (Cambridge University Press, Cambridge, 2106);
%arXiv:1501.02691}% end \textcolor{teal}{

%Page82,Stapp85,Tipler00,Bacciagaluppi02,TimpsonBrown02,Vaidman15

%\textcolor{orange}{\bibitem{Weissman99}M.~B.~Weissman, ``Emergent measure-dependent probabilities from modified quantum dynamics without state-vector reduction,''
%{\em Found. Phys. Lett.}\/ {\bf 12}  407-426   (1999); arXiv:quant-ph/9906127.
%}%end  \textcolor{orange}{ -- REMOVED
%
%\textcolor{orange}{
%\bibitem{Hanson03} 
%  R.~D.~Hanson, ``When worlds collide: Quantum probability from observer selection?,''
%{\em Found. Phys.} {\bf 33}, 1129-1150 (2003);  quant-ph/0108070. 
%}%end  \textcolor{orange}{ -- REMOVED
% 
%\textcolor{orange}{
%\bibitem{Hanson06}R.~D.~Hanson, ``Drift-diffusion in mangled worlds quantum mechanics,''
%{\em Proc. Roy. Soc. A}\/,, {\bf 462},  1619-1627 (2006); quant-ph/0303114.
%}%end  \textcolor{orange}{ -- REMOVED
%
%\bibitem{Wallace12}D.~Wallace,  {\em The Emergent Multiverse: Quantum Theory According to the Everett Interpretation}\/, (Oxford University Press, Oxford, 2012).
%
\bibitem{Kolmogorov56}A.~N.~Kolmogorov, {\em Foundations of the Theory of Probability}, 2nd English ed. (Chelsea Publishing Company, New York, 1956).


\bibitem{deFinetti89} B.~de~Finetti, ``Probabilism: A critical essay on the theory of probability and on the value of science,'' {\em Erkenntnis}\/
{\bf 31}, 169-223 (1989).% end textcolor{blue}{


\bibitem{Koopmans40}B.~O.~Koopmans, ``The axioms and algebra of intuitive probability,'' {\em Annals of Mathematics}\/, 2nd series, {\bf 41}, 269-292 (1940).

\bibitem{SEPprobability} A.~H\'{a}jek, ``Interpretations of Probability,'' {\em The Stanford Encyclopedia of Philosophy}\/ (Fall 2019 Edition), 
E.~ N.~Zalta ed., https://plato.stanford.edu/archives/fall2019/entries/probability-interpret/.

\bibitem{Deutsch16} D.~Deutsch, ``The logic of experimental tests, particularly of Everettian quantum theory,''  {\em Studies in History and Philosophy of Modern Physics}\/ {\bf 55} 24-33 (2016).%end \textcolor{teal}{


\bibitem{deFinetti17}  B.~de Finetti, {\em Theory of Probability: A Critical Introductory Treatment}\/ (John Wiley \& Sons Ltd., West Sussex, 2017).%end \textcolor{teal}{

%\begin{comment}
%
%\clearpage
%
%\mbox{}
%
%\clearpage
%
%
%\bibitem{Geroch84}R.~Geroch, ``The Everett interpretation,'' {\em No\^{u}s}\/ {\bf 18}, 617-633 (1984).
%
%
%\bibitem{Rubin10}M.~A.~Rubin, ``Explanation, evolution and subjective probability in Everett quantum mechanics with positive preclusion,'' arXiv:0912.4239. (2010). 
%
%\bibitem{Sorkin94} R.~D.~Sorkin, ``Quantum mechanics as quantum measure theory,'' {\em Mod. Phys. Lett.}\/ {\bf A9}  3119-3128 (1994); gr-qc/9401003.
%
%\bibitem{Sorkin97} R.~D.~Sorkin, ``Quantum measure theory and its interpretation,'' in  D.~H.~Feng and B.-L.~Hu, eds., 
%{\em Quantum Classical Correspondence: Proceedings of the 4th Drexel Symposium on Quantum Nonintegrability, held Philadelphia, September 8-11, 1994}\/, 229-251 (International Press, Cambridge Mass., 1997); gr-qc/9507057.
%
%\bibitem{Sorkin07a} R.~D.~Sorkin, ``An exercise in `anhomomorphic logic' ,''
%{\em J. Phys.}\/ Conf. Ser. {\bf 67}, 012018 (2007); quant-ph/0703276.
%
%\bibitem{Sorkin07b} R.~D.~Sorkin, ``Quantum dynamics without the wave function,'' {\em J. Phys. A} {\bf 40}, 3207-3222 (2007); quant-ph/0610204.
%
%\bibitem{DowkerGhaziTabatabai08b} F.~Dowker and Y.~Ghazi-Tabatabai, ``Dynamical wave function collapse models in quantum measure theory,''
% {\em J. Phys. A}\/ {\bf 41}, 205306 (2008); arXiv:0712.2924.
%
%\bibitem{DowkerGhaziTabatabai08a} F.~Dowker and Y.~Ghazi-Tabatabai, ``The Kochen-Specker theorem revisited in quantum measure theory,'' 
%{\em J. Phys. A}\/ {\bf  41}, 105301 (2008); arXiv:0711.0894.
%
%\bibitem{Barnettetal07}M.~Barnett, F.~Dowker and D.~Rideout, ``Popescu-Rohrlich boxes in quantum measure theory,''
%  {\em J. Phys. A} {\bf 40},  7255-7264 (2007); quant-ph/0605253.
%
%\bibitem{Craigetal07}D.~Craig, F.~Dowker, J.~Henson, S.~Major, D.~Rideout and R.~D.~Sorkin, ``A Bell inequality analog in quantum measure theory,''
%{\em  J. Phys. A}\/ {\bf 40}, 501-523 (2007);  quant-ph/0605008.
%
%
%\bibitem{Ghazi-TabatabaiWallden09a}Y.~Ghazi-Tabatabai and P.~Wallden,
%``Dynamics \& predictions in the co-event interpretation,''
%     {\em J. Phys. A}\/, {\bf 42}, 235303 (2009); arXiv:0901.3675.
%
%\bibitem{Ghazi-Tabatabai09}Y.~Ghazi-Tabatabai, ``Quantum measure theory: A new interpretation,''arXiv:0906.029.
% 
%\bibitem{Ghazi-TabatabaiWallden09b}Y.~Ghazi-Tabatabai and P.~Wallden,
% ``The emergence of probabilities in anhomomorphic logic,''
% {\em J. Phys.}\/ Conf. Ser. {\bf 174}, 012054 (2009);  arXiv:0907.0754. 
% 
%\bibitem{Gudder09} S.~Gudder, ``An anholonomic logic for quantum mechanics,'' arXiv:0910.3253 (2009).
%	
%\bibitem{Gudder11} S.~Gudder, ``Quantum measures and the coevent interpretation,'' {\em Repts. Math.
%Phys}\/  {\bf 67}, 137-156 (2011); arXiv:1005.2242.
% 
% \bibitem{FraucaSorkin17} A.~M.~Frauca, R.~D.~Sorkin, ``How to measure the quantum measure,'' {\em Int. J. Theor. Phys.}\/ 
% {\bf 56}, 232?258 (2017);arXiv:1610.02087  
% 
%
%
%\bibitem{Galvan07a} B.~Galvan, ``Typicality vs. probability in trajectory-based formulations of quantum mechanics,''
%{\em Found. Phys.} {\bf 37}, 1540-1562 (2007); arXiv:quant-ph/0605162.
%
%\bibitem{Galvan07b} B.~Galvan, ``Origin of which-way information and generalization of the Born rule,'' 
%arXiv:0705.2877.
%
%
%\bibitem{Galvan08a} B.~Galvan,  ``Quantum mechanics and imprecise probability,''  {\em J. Stat. Phys.}\/ {\bf 131}, 1155-1167
%(2008); arXiv:0711.3000.
%
%
%
%\bibitem{Galvan08b} B.~Galvan, ``Generalization of the Born rule,''
%{\em  Phys. Rev. A}\/ {\bf 78}, 042113 (2008); arXiv:0806.4935.
%
%
%
%
%
%
%
%\bibitem{Cournot1843} A.~A.~Cournot, {\em Exposition de la th\'{e}orie des chances et des probabilit\'{e}s}\/  (Hachette,
%Paris, 1843). Reprinted as Volume I of B.~Bru ed., {\em Antoine-Augustin Cournot, Oeuvres
%compl\`{e}tes}\/ (Vrin, Paris, 1984).
%
%
%\bibitem{ShaferVovk06}G.~Shafer and V.~Vovk, ``The sources of Kolmogorov's {\em Grundbegriffe},'' {\em Statistical Science} {\bf 21},
%70-98 (2006).
%
%
%
%
%\bibitem{Kolmogorov56}A.~N.~Kolmogorov, {\em Foundations of the Theory of Probability}, 2nd English ed. (Chelsea Publishing Company, New York, 1956).
%
%
%
%\bibitem{Rubin03}M.~A.~Rubin, ``Relative frequency and probability in the Everett interpretation of Heisenberg-picture quantum mechanics,'' {\em Found. Phys.} {\bf 33}  379-405 (2003); 
%quant-ph/0209055.
%
%\bibitem{dEspagnat76}  B.~d'Espagnat, {\em Conceptual Foundations of Quantum Mechanics}\/,
%2nd edn. (W.~A.~Benjamin, Reading, MA, 1976).
%
%\bibitem{Rubin01}  M.~A.~Rubin,  ``Locality in the Everett interpretation of 
%Heisenberg-picture quantum mechanics,'' {\em Found. Phys. Lett.}\/ {\bf 14}, 301-322 (2001);
%quant-ph/0103079.
%
%\bibitem{Rubin04}M.~A.~Rubin, ``There is no basis ambiguity in Everett quantum mechanics,'' 
%{\em Found. Phys. Lett.}\/ {\bf 17}, 323-341 (2004); arXiv:quant-ph/0310186.
%
%\bibitem{Rubin06}M.~A.~Rubin, ``Spatial degrees of freedom in  Everett quantum mechanics,'' {\em Found. Phys.}\/ {\bf 36}, 1115-1159 (2006);arXiv:quant-ph/0511188.
%
%\bibitem{Beckenstein01} J.~D.~Beckenstein, ``Limitations on quantum information
%from black hole physics,'' {\em Acta Phys.Polon.}\/ {\bf B32}, 3555-3570 
%(2001); quant-ph/0110005.
%
%
%\bibitem{Deutsch02} D.~Deutsch, ``The structure of the multiverse,'' {\em Proc. Roy. Soc. London Ser. A}\/
%{\bf 458} 2911-2923 (2002); arXiv:quant-ph/0104033.
%
%\bibitem{Stein84}H.~Stein, ``The Everett interpretation of quantum mechanics: Many worlds or none?,''
%{\em No\^{u}s}\/ {\bf 18}. 635-652 (1984).
%
%\bibitem{Farhietal89}E.~Farhi, J.~Goldstone and S.~Gutmann, ``How probability arises in quantum mechanics,'' {\em Ann. Phys.  (NY)}\/ {\bf 192},
%368-382 (1989).
%
%\bibitem{Renyi70} A.~R\'{e}nyi, {\em Foundations of Probability}\/ (Holden-Day, Inc., San Francisco, 1970).
%
%
%\bibitem{Oldhametal09} K.~Oldham, J.~Myland and J.~Spanier, {\em An Atlas of Functions}\/, 2nd ed.(Springer Science+Business Media, LLC, New York, 2009).
%
%
%
%\bibitem{deFinetti89} B.~de~Finetti, ``Probabilism: A critical essay on the theory of probability and on the value of science,'' {\em Erkenntnis}\/
%{\bf 31}, 169-223 (1989).
%
%\bibitem{Lewis80} D.~Lewis, ``A  subjectivist's guide to objective chance,'' in R.~Jeffrey, ed., 
%{\em Studies in Inductive Logic and Probability, Vol. II.}\/ (University of California Press, Berkeley, 1980).
%
%
%\bibitem{BrownBenPorath20}H.~R.~Brown and G.~Ben~Porath, ``Everettian probabilities, the Deutsch-Wallace theorem and the Principal Principle,''
%in
%M.~Hemmo  and O.~Shenker, eds.,
%{\em Quantum, Probability, Logic:
%The Work and Influence of Itamar Pitowsky}\/
%(Springer Nature Switzerland AG, Cham, Switzerland, 2020)
%
%\bibitem{Earman18}J.~Earman, ``The Relation between credence and chance: Lewis' `Principal Principle' is a theorem of quantum probability theory,'' http://philsci-archive.pitt.edu/id/eprint/14822 (2018). 
%
%\bibitem{Davies74}P.~C.~W.~Davies, {\em The Physics of Time Asymmetry} (University of California Press, Berkeley, 1974).
%
%\bibitem{Teglasetal07}E.~T\'{e}glas, V.~Girotto, M.~Gonzalez and L.~Bonatti, ``Intuitions of probabilities shape expectations about the future at 12 months and beyond,'' {\em PNAS} \/{\bf 104}19156-19159 (2007).
%
%
%\bibitem{TeglasBonatti16}E.~T\'{e}glas and L.~Bonatti, ``Infants anticipate probabilistic but not deterministic outcomes,'' {\em Cognition}\/, {\bf 157} 227-236 (2016).
%
%\bibitem{XuGarcia08}F.~Xu and V.~Garcia, ``Intuitive statistics by 8-month-old infants,'' {\em PNAS}\/ {\bf 105}, 5012-5025 (2008).
%
%\bibitem{HanusCall14}D.~Hanus and J.~Call ``When maths trumps logic: probabilistic judgements in chimpanzees,'' {\em Biology Letters}\/
%{\bf 10},  20140892 (2014).
%
%\bibitem{Rakoczyetal14} H.~Rakoczy, A.~Cl\"{u}ver, L.~Sauke, N.~Stoffregen, A.~Gr\"{a}bener, J.~Migura and J.~Call, ``Apes are
%intuitive statisticians,'' {\em Cognition} {\bf 131} 60-68 (2014).
%
%\bibitem{Tecwynetal17} E.~C.~Tecwyn, S.~Denison, E.~J.~E.~Messer and D.~Buchsbaum, ``Intuitive probabilistic inference in capuchin monkeys,'' {\em Anim. Cogn.}\/ {\bf 20} 243-256 (2017).
%
%\bibitem{DePetrilloRosati19}F.~De~Petrillo and A.~G.~Rosati, ``Rhesus macaques use probabilities to predict future events,'' 
%{\em Evolution and Human Behavior}\/ {\bf 40} 436-446 (2019).
%
%
%
%
%
%
%
%
%
%\bibitem{DenisonXu19}S.~Denison and F.~Xu, ``Infant statisticians: The origins of reasoning under uncertainty,'' {\em Perspectives on Psychological Science}\/ {\bf 14}, 499-509 (2019).
%
%
%\bibitem{deFinetti17} B.~de Finetti, {\em Theory of Probability: A Critical Introductory Treatment}\/ (John Wiley \& Sons Ltd., West Sussex, 2017)
%
%
%\bibitem{Forrester07} A.~Forrester, ``Decision theory and information propagation in quantum physics,'' {\em History and Philosophy of Science Part B: Studies In History and Philosophy of Modern Physics}\/ {\bf 38}, 815-831 (2007); arXiv:quant-ph/0604133.
%%
%%\bibitem{SebensCarroll18}C.~T.~Sebens and S.~M.~Carroll, ``Self-locating uncertainty and the
%%origin of probability in Everettian quantum mechanics,''
%%{\em Brit. J. Phil. Sci.}\/ {\bf 69},  25-74 (2018).
% 
%
%
%\bibitem{DeutschHayden00} D.~Deutsch and P.~Hayden, ``Information flow in entangled quantum systems,'' {\em Proc. R. Soc. Lond.}\/ {\bf A.456}, 1759-1774 (2000);arXiv:quant-ph/9906007. 
%
%\bibitem{Rubin02}M.~A.~Rubin, ``Locality in the Everett Interpretation of Quantum Field Theory,''
%{\em Found.Phys.} {bf 32}, 1495-1523 (2002); arXiv:quant-ph/0204024.
%
%
%\bibitem{HewittHorsmanVedral07a}C.~Hewitt-Horsman and V.~Vedral, ``Developing the Deutsch-Hayden approach to quantum mechanics,'' {\em New Journal of Physics}\/ {\bf 9} 135 (2007); arXiv:quant-ph/0609085.
%
%\bibitem{HewittHorsmanVedral07b} C.~Hewitt-Horsman and  V.~Vedral, ``Entanglement without nonlocality,'' {\em Phys. Rev.}\/  {\bf A 76}, 062319 (2007); arXiv:quant-ph/0611237.
%
%\bibitem{Rubin11}M.~A.~Rubin, ``Observers and locality in Everett quantum field theory,'' 
%{\em Found.Phys.}\/ {\bf 41}, 1236-1262 (2011);arXiv:0909.2673.
%
%\bibitem{Deutsch12} D.~Deutsch, ``Vindication of quantum locality,'' {\em Proc. R. Soc. A}\/ {\bf 468}, 531-544 (2012); arXiv:1109.6223.
%
%\bibitem{Bedard20}C.~A.~B\'{e}dard, ``Topics on quantum locality," arXiv:2001.09646 (2020).
%
%\bibitem{KuypersDeutsch20}S.~Kuypers and D.~Deutsch, ``Everettian relative states in the Heisenberg picture,'' arXiv:2008.02328 (2020). 
%
%\bibitem{Heisenberg58}W.~Heisenberg,  {\em Physics and Philosophy}\/ (Prometheus Books, Amherst, 1958).
%
%\bibitem{Heisenberg74}W.~Heisenberg, {\em Across the Frontiers}\/, (Harper and Row, Publishers, Inc., New York 1974).
%
%\bibitem{Fleming92}G.~Fleming, ``The actualizion of potentialities in contemporary quantum theory,'' {\em The Journal of Speculative Philosophy}\/ {\bf 6} 259-276 (1992).
%
%\bibitem{Pangle14}T.~Pangle, ``On Heisenberg's key statement concerning ontology,'' {\em The Review of Metaphysics}\/ {\bf 67}, 835-859 (2014).
%
%\bibitem{Bschir16}K.~Bschir, ``Potentiality in natural philosophy,'' in T.~E.~Eastman, M.~Epperson and D.~R.~Griffin, eds.,  
%{\em Physics and Speculative Philosophy: Potentiality in Modern Science}\/ (De Gruyter, Berlin, 2016).
%
%\bibitem{Jaeger17}G.~Jaeger, ``Quantum potentiality revisited,'' {\em Phil. Trans. R. Soc. A}\/ {\bf 375}, 20160390 (2017).
%
%\bibitem{Jaeger17b}G.~Jaeger, `` `Wave-packet reduction' and the quantum character of the actualization of potentia,'' {\em Entropy}\/
%{\bf 19}, 513 (2017).
%
%\bibitem{BishopBrenner17}R.~R.~Bishop and J.~E.~Brenner, ``Potentiality, actuality and non-separability in quantum and classical physics: Res potentiae in the macroscopic world,'' arXiv:1801.01471 (2017).
%
%\bibitem{Kastner18} R.~E.~Kastner, ``Taking Heisenberg's potentia seriously,'' {\em Int. J. Quant. Found.}\/ {\bf 4}, 158-172 (2018).
%
%\bibitem{Kistler18}M.~Kistler, ``Potentiality in physics,'' in K.~Engelhard and M.~Quante, eds.,  {\em Potentiality in Physics}\/
%(Springer, Dordrech, 2018).
%
%\bibitem{Sanders18}G.~Sanders, ``An Aristotelian approach to quantum mechanics,'' https://www.academia.edu/download/56222749/An\_Aristotelian\_Approach\_to\_ Quantum\_Mechanics.pdf  (2018).
%
%
%\bibitem{DeWitt98}B.~DeWitt, ``The quantum mechanics of isolated systems,'' {\em Int. J. Mod. Phys. A}\/ {\bf 13}, 1881-1916 (1998).
%
%\bibitem{Polley00} L.~Polley, ``Quantum-mechanical probability from the symmetries of two-state systems,''   quant-ph/9906124 (2000).
%
%\bibitem{Wallace02}D.~Wallace, ``Quantum probability and decision theory, revisited,'' arXiv:quant-ph/0211104 (2002).
%
%\bibitem{Wallace03} D.~Wallace, ``Everettian rationality: defending Deutsch's approach
%to probability in the Everett interpretation,'' {\em Stud. 
%Hist. Phil. Mod. Phys.}\/ {\bf 34}, 415-439 (2003); quant-ph/0303050.
%
%\bibitem{Wallace07}D.~Wallace, ``Quantum probability from subjective likelihood: Improving
%on Deutsch's proof of the probability rule,''{\em Stud. 
%Hist. Phil. Mod. Phys.}\/  {\bf 38}, 311-332 (2007); quant-ph/0312157.
%
%
%
%\
%\bibitem{VanEsch07}P.~Van ~Esch, ``On the Born rule and the Everett programme,''
%{\em Annales de la Fondation Louis de Broglie}\/ {\bf 32}, 51-59  (2007),
%
%\bibitem{Barrett17}J.~A.~Barrett, ``Typical worlds,'' {\em Studies in History and Philosophy of Science Part B: Studies in History and Philosophy of Modern Physics}\/
%{\bf 58}, 31-40 (2017); arXiv:1912.05312
%
%\bibitem{Vaidman20}L.~Vaidman,
%``Derivations of the Born rule,'' in
%M.~Hemmo  and O.~Shenker, eds.,
%{\em Quantum, Probability, Logic:
%The Work and Influence of Itamar Pitowsky}\/
%(Springer Nature Switzerland AG, Cham, Switzerland, 2020)
%
%\bibitem{Page82}D.~N.~Page, ``The Einstein-Podolsky-Rosen physical reality is completely described by quantum mechanics,''  {\em Phys. Lett.} {\bf  A91}, 57-60 (1982).
%
%\bibitem{Stapp85} H.~P.~Stapp, ``Bell's theorem and the foundations of quantum physics,''
%{\em Am. J. Phys.}\/ {\bf 53}, 306 (1985).
%
%\bibitem{Tipler00} F.~J.~Tipler, ``Does quantum nonlocality exist? Bell's theorem and the many-worlds interpretation,'' arXiv:quant-ph/0003146 (2000).
%
%\bibitem{Bacciagaluppi02} G.~Bacciagaluppi, ``Remarks on space-time and locality in Everett's interpretation.,'' in  
%T.~Placek and J.~Butterfield, {\em Non-locality and Modality}\/ (Kluwer Academic Publishers, Dordrecht, 2002); philsci-archive.pitt.edu/id/eprint/504. 
%
%\bibitem{TimpsonBrown02} C.~G.~Timpson and  H.~R.~Brown, ``Entanglement and relativity,'' in   R.~Lupacchini and V.~Fano,  eds., {\em Understanding Physical Knowledge}\/ (CLUEB, Bologna, 2002).
%
%\bibitem{Vaidman15} L.~Vaidman, ``The Bell inequality and the many-worlds interpretation,'' in M.~Bell and
%S.~Gao, eds., {\em Quantum Nonlocality and Reality: 
%50 Years of Bell's Theorem}\/ (Cambridge University Press, Cambridge, 2106);
%arXiv:1501.02691
%
%\bibitem{Bell04} J.~S.~Bell, {\em Speakable and Unspeakable in Quantum Mechanics}\/, 2nd ed. (Cambridge University Press, Cambridge, 2004).
%
%\bibitem{Bell64}J.~S~Bell, ``On the Einstein-Podolsky-Rosen paradox,'' reprinted in \cite{Bell04} (1964).
%
%\bibitem{Bell75}J.~S~Bell, ``The theory of local beables,'' reprinted in \cite{Bell04} (1975). 
%
%\bibitem{Bell81}J.~S~Bell,  ``Bertlmann's socks and the nature of reality,'' reprinted in \cite{Bell04} (1981).
%
%\bibitem{Bell90}J.~S~Bell, ``La nouvelle cuisine,'' reprinted in \cite{Bell04} (1990).
%
%\bibitem{Norsen06}T.~Norsen, ``Bell locality and the nonlocal character of nature,''
%{\em Found. Phys. Lett.} {\bf 19}, 633-655 (2006); arXiv:quant-ph/0601205.
%
%
%\bibitem{Blaylock10}G.~Blaylock, ``The EPR paradox, Bell's inequality, and the question of locality,''
%{\em Am. J. Phys.}\/ {\bf 78}, 111-120 (2010);arXiv:0902.3827.
%
%\bibitem{Maudlin14} T.~Maudlin, ``What Bell did,''  {\em J. Phys. A: Math. Theor.}\/ {\bf 47} 424010 (2014); arXiv:1408.1826.
%
%\bibitem{Myrvoldetal20} W.~Myrvold, M.~Genovese and A.~Shimony, ``Bell's Theorem,'' {\em The Stanford Encyclopedia of Philosophy }\/(Fall 2020 Edition), Edward N. Zalta ed., https://plato.stanford.edu/archives/fall2020/entries/bell-theorem/.
%
%\bibitem{Deutsch16} D.~Deutsch, ``The logic of experimental tests, particularly of Everettian quantum theory,''  {\em Studies in History and Philosophy of Modern Physics}\/ {\bf 55} 24-33 (2016).
%
%\bibitem{Zurek03} W.~H.~Zurek, ``Environment-assisted invariance, entanglement, and probabilities in quantum physics,'' 
%{\em Phys. Rev. Lett.}\/ {\bf 90} 120404 (2003); arXiv:quant-ph/0211037.
%
%\bibitem{Zurek05} W.~H.~Zurek, ``Probabilities from entanglement, Born's rule 
%$p_k=|\psi_k|^2$\/ from envariance,'' {\em Phys. Rev. A}\/ {\bf 71} 052105  (2005); arXiv:quant-ph/0405161.
%
%\bibitem{Vaidman98} L.~Vaidman,  ``On schizophrenic experiences of the neutron or why we should believe in the many-worlds interpretation of quantum theory,''  {\em International Studies in the Philosophy of Science}\/ {\bf 12} 245-261 (1998); arXiv:quant-ph/9609006.
%
%
%\bibitem{McQueenVaidman19} K.~J.~McQueen and L.~Vaidman, ``In defence of the self-location uncertainty account of probability in the many-worlds interpretation,'' {\em Studies in History and Philosophy of Science Part B: Studies in History and Philosophy of Modern Physics}\/
%{\bf 66} 14-23 (2019); 
%
%\bibitem{Kent15}A.~Kent, ``Does it make sense to speak of self-locating uncertainty in the universal wave function? Remarks on Sebens and Carroll,'' {\em Foundations of Physics} {\bf 45} 211?217 (2015); arXiv:1408.1944.
%%
%%\bibitem{SebensCarroll18}C.~T.~Sebens and S.~M.~Carroll, ``Self-locating uncertainty and the
%%origin of probability in Everettian
%%quantum mechanics,''
%%{\em Brit. J. Phil. Sci.}\/ {\bf 69} 25?74 (2018);arXiv:1405.7577 
%
%
%\end{comment}
%
%
%

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{wider}
\usepackage{xcolor}
\usepackage{soul}
\usepackage{comment}
\usepackage{cite}
\usepackage{fnpct}

\begin{document}

\newcommand{\sqvb}{\ensuremath{ \langle \!\langle 0 |} }
\newcommand{\sqvk}{\ensuremath{ | 0 \rangle \!\rangle } }
\newcommand{\sqvn}{\ensuremath{ \langle \! \langle 0 |  0 \rangle \! \rangle} }
\newcommand{\wh}{\ensuremath{\widehat}}
\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}
\newcommand{\ra}{\ensuremath{\rangle}}
\newcommand{\la}{\ensuremath{\langle}}
\newcommand{\rra}{\ensuremath{ \rangle \! \rangle }}
\newcommand{\lla}{\ensuremath{ \langle \! \langle }}
\newcommand{\str}{\rule[-.125cm]{0cm}{.5cm}}
\newcommand{\pr}{\ensuremath{^{\;\prime}}}
\newcommand{\ppr}{\ensuremath{^{\;\prime \prime}}}
\newcommand{\da}{\ensuremath{^\dag}}
\newcommand{\as}{^\ast}
\newcommand{\eps}{\ensuremath{\epsilon}}
\newcommand{\ve}{\ensuremath{\vec}}
\newcommand{\ka}{\kappa}
\newcommand{\non}{\ensuremath{\nonumber}}
\newcommand{\lf}{\ensuremath{\left}}
\newcommand{\rt}{\ensuremath{\right}}
\newcommand{\al}{\ensuremath{\alpha}}
\newcommand{\dfn}{\ensuremath{\equiv}}
\newcommand{\ga}{\ensuremath{\gamma}}
\newcommand{\ti}{\ensuremath{\tilde}}
\newcommand{\wti}{\ensuremath{\widetilde}}
\newcommand{\hs}{\ensuremath{\hspace*{.5cm}}}
\newcommand{\bet}{\ensuremath{\beta}}
\newcommand{\om}{\ensuremath{\omega}}
\newcommand{\kp}{\ensuremath{\kappa}}

\newcommand{\cO}{\ensuremath{{\cal O}}}
\newcommand{\cS}{\ensuremath{{\cal S}}}
\newcommand{\cF}{\ensuremath{{\cal F}}}
\newcommand{\cX}{\ensuremath{{\cal X}}}
\newcommand{\cZ}{\ensuremath{{\cal Z}}}
\newcommand{\cG}{\ensuremath{{\cal G}}}
\newcommand{\cR}{\ensuremath{{\cal R}}}
\newcommand{\cV}{\ensuremath{{\cal V}}}
\newcommand{\cC}{\ensuremath{{\cal C}}}
\newcommand{\cP}{\ensuremath{{\cal P}}}
\newcommand{\cH}{\ensuremath{{\cal H}}}
\newcommand{\cN}{\ensuremath{{\cal N}}}
\newcommand{\cE}{\ensuremath{{\cal E}}}

\newcommand{\pup}{\ensuremath{^{(p)}}}
\newcommand{\prpr}{\ensuremath{\prime \prime }}

\newcommand{\hsp}{\ensuremath{\hspace*{5mm} }}
\newcommand{\sbp}{\ensuremath{_{[p]} }}

\newcommand{\erf}{\mathop{\mathrm{erf}}}

\newcommand{\xxx}[1]{}
\newcommand{\yyy}[1]{}
\newcommand{\zzz}[1]{}
\newcommand{\xyz}[1]{}

\title{\bf Reverse quantum speed limit and minimum Hilbert space norm}

\author{
Mark A. Rubin\\
\mbox{}\\
markallenrubin@yahoo.com\\
}
\date{\mbox{}}
\maketitle
\begin{abstract}
The reverse quantum speed limit (RQSL)  gives an upper limit to the time for evolution between initial and final quantum states.  We show that, in conjunction with the existence of a minimum time scale, the RQSL implies a lower limit to the norm of the change in a quantum state, and confirm that this limit is satisfied in two-state and ideal-measurement models.  Such a lower limit is of  relevance for interpretational issues in probability and for understanding the meaning of probability in Everett quantum theory.


\mbox{}

\noindent Key words: reverse quantum speed limit, minimum Hilbert space norm, minimum time scale,  discrete Hilbert space, probability, preclusion,  Everett interpretation

\end{abstract}

\section{Introduction}\label{SecIntro}
\xxx{SecIntro}

The spacetime that provides the mathematical framework within which both classical and quantum fields are defined is a continuum. Nevertheless, it has been argued for decades that  the combination of quantum mechanics and general relativity implies the existence of a minimum length scale at or near the  Planck
 length.\footnote{A  device-independent argument for such a minimum length scale is presented in
  \cite{Calmetetal04}.
  For a review of this subject, see \cite{Hossenfelder13}.}
Relativistic covariance then implies the existence of a  corresponding minimum time scale comparable in magnitude to the Planck time \cite{BHZ05, FaizalKhalilDas16}.

Buniy, Hsu and Zee \cite{BHZ05}  have argued that the existence of a minimum length scale implies a minimum scale to the norms of states in quantum-mechanical Hilbert space,  such that components of the wavefunction with norm below this scale can be removed from the wavefunction.\footnote{Recently Calmet and Hsu \cite{CalmetHsu21} have presented an argument for such a Hilbert-space minimum norm based on a minimum scale to angular measurement coming from quantum mechanics and general relativity.}  The existence of such a minimum Hilbert-space norm 
provides a  solution to the problem of ``maverick'' states\cite{DeWitt70} in Everett quantum theory\cite{Everett57, DeWittGraham73, Price95, Barrett99, Vaidman02,  HewittHorsman09, Saundersetal10, Wallace12, Wallace13} and thus allows for an understanding of the origin and nature of probability in that theory  \cite{BHZ06, Hsu12, Hsu17, Hsu21, Rubin21}.

%

Hsu\cite{Hsu21} 
has presented an estimate for the norm of the change in a quantum state in the presence of a minimum time scale. In the present note we employ the reverse quantum speed limit (RQSL) of Mohan and Pati \cite{MohanPati20} to derive a lower limit to this norm.

\section{Reverse quantum speed limit}\label{SecRQSL}
\xxx{SecRQSL}

The RQSL gives an upper limit to the time $T$\/ for an initial state $|\psi(0)\ra$\/ to evolve to a final state $|\psi(T)\ra$\/.  
The expression for this limit involves the reference section $|\chi(t)\ra$\/, a function of the state vector $|\psi(t)\ra$\/ and the initial time (here taken to be $t=0$\/) defined as
\be
|\chi(t)\ra = \frac{\la \psi(t) | \psi(0) \ra }{| \la \psi(t) | \psi(0) \ra |} |\psi(t)\ra. \label{refsecdef}
\ee
\xxx{refsecdef}
\yyy{DHS-6, last box}
The length of the reference section for evolution from time $t=0$\/ to time $t=T$\/ is defined to be
\be
l(\chi(t))|_{0}^{T} =\int_{0}^{T} \la\dot{\chi}(t)|\dot{\chi}(t)\ra^{\frac{1}{2}} dt. \label{refseclendef}
\ee
\xxx{refseclendef}
\yyy{p. DHS-6, box 2}
In terms of the length (\ref{refseclendef}), the RQSL  is
\be 
T \leq \frac{\hbar l(\chi(t))|_{0}^{T} }{\Delta H},\label{RQSL}
\ee
\xxx{RQSL}
\yyy{p. DHS-6, box 1}
where $\Delta H$\/ is the square root of the expectation value of the variance of the Hamiltonian $\wh{H}$\/ in the state $|\psi(t)\ra$\/:
\be
\Delta H=\left(\la \psi(t)|\wh{H}^2|\psi(t)\ra-\la \psi(t)|\wh{H}|\psi(t)\ra^2\right)^\frac{1}{2}.\label{DeltaHdef}
\ee
\xxx{DeltaHdef}
\yyy{p. DHS-16, box 2}
Here we will only consider the case of time-independent Hamiltonians,\footnote{For the generalization to time-dependent Hamiltonians, see 
\cite{MohanPati20}.}
in which case $\Delta H$\/ is time-independent and the formula (\ref{DeltaHdef}) can be evaluated at any convenient time, in particular at $t=0$\/.

\section{Minimum Hilbert space norm}

Consider the time evolution to occur over a time interval $0 \leq t \leq \Delta t$\/ short enough so that we can approximate the derivative of the reference section (\ref{refsecdef}) during this interval as 
\be
|\dot{\chi}(t)\ra = \frac{|\chi(\Delta t)\ra -|\chi(0)\ra }{\Delta t}. \label{chiapprox}
\ee
\xxx{chiapprox}
\yyy{p. DHS-6, box 3}
Clearly the range of applicability of subsequent results will depend on the range of validity of the approximation (\ref{chiapprox}); we will return to this issue below.  

Using (\ref{chiapprox}) in (\ref{refseclendef}) with $T=\Delta t$\/,
\bea
l(\chi(t))|_{0}^{\Delta t}&=&\left[\left(\frac{\la\chi(\Delta t)| -\la\chi(0)| }{\Delta t}\right)\left(\frac{|\chi(\Delta t)\ra -|\chi(0)\ra }{\Delta t}\right)\right]^\frac{1}{2}\Delta t \nonumber \\
&=&\left[\left(\la\chi(\Delta t)| -\la\chi(0)| \right)\left(|\chi(\Delta t)\ra -|\chi(0)\ra \right)\right]^\frac{1}{2} \label{refseclen2}
\eea
\xxx{refseclen2}
\yyy{p. DHS-6, box 5}
Using (\ref{refsecdef}) in (\ref{refseclen2}), along with the normalization of $|\psi(t)\ra$\/,  yields  
\be
l(\chi(t))|_{0}^{\Delta t} =  \left[\la\psi(\Delta t)|\psi(\Delta t)\ra  -2|\la\psi(\Delta t)|\psi(0)\ra| +1\right]^\frac{1}{2}\label{refseclen3}
\ee
\xxx{refseclen3}
\yyy{p. DHS-8}
Expressing $|\psi(\Delta t )\ra$\/ in terms of $|\Delta \psi (\Delta t)\ra$\/, the change in $|\psi(t)\ra$\/, i.e. 
\be
|\psi(\Delta  t)\ra=|\psi(0)\ra + |\Delta \psi (\Delta t) \ra, \label{psitodeltapsi}
\ee
\xxx{psitodeltapsi}
\yyy{p. DHS-9, box 2}
and defining
\be
z=1+\la\Delta \psi (\Delta t) |\psi(0)\ra \label{zdef}
\ee
\xxx{zdef}
\yyy{p. DHS-12, box 4, p. DHS-31 box 3}
we obtain from (\ref{refseclen3})
\be
\left(l(\chi(t))|_{0}^{\Delta t}\right)^2 = \la \Delta \psi (\Delta t)| \Delta \psi (\Delta t)\ra + 2\left({\rm Re}(z) -|z|\right). \label{refseclenitoz}
\ee
\xxx{refseclenitoz}
\yyy{p. DHS-31, last box}
From the RQSL, eq. (\ref{RQSL}),
\be
\left(l(\chi(t))|_{0}^{\Delta t}\right)^2 \ge \left(\frac{\Delta t \Delta H}{ \hbar}\right)^2,\label{RSQL2}
\ee
\xxx{RSQL2}
\yyy{p. DHS-31, next-to-last box}
so, with (\ref{refseclenitoz}), we have
\be
\la\Delta \psi (\Delta t) | \Delta \psi (\Delta t) \ra -\left(\frac{\Delta t \Delta H}{ \hbar}\right)^2 \ge 2\left(|z|-{\rm Re}(z)\right). \label{almostthere}
\ee
\xxx{almostthere}
\yyy{p. DHS-31, box 1}
But for any complex number $z$\/
\be
|z|-{\rm Re}(z) \ge 0, \label{complexnumbers}
\ee
\xxx{complexnumbers}
\yyy{p. DHS-33}
so 
\be
\la \Delta \psi (\Delta t)| \Delta \psi(\Delta t) \ra^\frac{1}{2} \ge \frac{\Delta t \Delta H}{\hbar}. \label{normlimit}
\ee
\xxx{normlimit}
\yyy{ p. DHS-12, last box}

Now suppose that time is in fact discrete; i.e., there is a lower limit $\Delta t_{min}$\/ during which time evolution can take place. Then
\be
\la \Delta \psi (\Delta t_{min})| \Delta \psi (\Delta t_{min})\ra^\frac{1}{2} \ge {\rm NormLim_{RQSL}},
\label{normlimit2}
\ee
\xxx{normlimit2}
where
\be
{\rm NormLim_{RQSL}} = \frac{\Delta t_{min} \Delta H}{\hbar}. \label{NLRQSLdef}
\ee
\xxx{NLRQSLdef}
That is, 
there is no Hamiltonian with expected variance $\left(\Delta H\right)^2$\/ in  state $|\psi(0)\ra$\/ which will in a time $\Delta t_{min}$\/ evolve $|\psi(0)\ra$\/ into a state 
$|\psi(\Delta t_{min})\ra$\/ closer in norm to $|\psi(0)\ra$\/ than ${\rm NormLim_{RQSL}}$\/.\footnote{Note that the estimate given by Hsu \cite[eq. (6)]{Hsu21} for the norm of the change in the state, which in our notation is  $\Delta t_{min} \la\psi(0)|\wh{H}^2|\psi(0)\ra^\frac{1}{2}/\hbar$\/, will always be greater than or equal to   ${\rm NormLim_{RQSL}}$\/.  Hsu argues that by considering the estimates for all possible initial states in a finite-dimensional Hilbert space one may arrive at a lower bound on the norm.  The minimum norm given in the present paper applies to  all states with a specified value of $\Delta H$\/.}

\section{Examples}

Since time appears to us to be continuous rather than discrete, we will take the approximation (\ref{chiapprox}) and the conclusion (\ref{normlimit2}) that follows from it be valid provided $\Delta t_{min}$\/ is much smaller than the smallest time scale characterizing changes in  $|\psi(t)\ra$\/.  We consider two examples:

\subsection{Two-state system}\label{SubsecTwoState}
\xxx{SubsecTwoState}

The Hilbert space $\cH_2$\/ is spanned by two vectors, $|\cS_1\ra$\/ and  $|\cS_2\ra$\/, which we take to be normalized eigenstates of the Hamiltonian $\wh{H}_{\cH_2}$\/:
\be
\wh{H}_{\cH_2}|\cS_i\ra=E_i|\cS_i\ra,\hspace*{1cm} \la \cS_i|\cS_j\ra=\delta_{i,j},\hspace*{1cm}i,j=1,2.\label{Hilbertspace2}
\ee
\xxx{Hilbertspace2}
\yyy{p. DHS-15, boxes 2,4,5}
The normalized initial state vector is
\be
|\psi_{\cH_2}(0)\ra=c_1|\cS_1\ra+c_2|\cS_2\ra,\hspace*{1cm}|c_1|^2+|c_2|^2=1.\label{psi0HS2}
\ee
\xxx{psi0HS2}
\yyy{p. DHS-15, box 1}
The lower limit to the norm of the change in the state vector from $t=0$\/ to $t=\Delta t_{min}$\/ is, from (\ref{DeltaHdef}), (\ref{NLRQSLdef}), (\ref{Hilbertspace2}) and (\ref{psi0HS2}),
\be
{\rm NormLim}_{{\rm RQSL} - \cH_2}=\left[|c_1|^2E_1^2 + |c_2|^2E_2^2-\left(|c_1|^2E_1+|c_2|^2E_2\right)^2\right]^\frac{1}{2}\frac{\Delta t_{min}}{\hbar}. 
\label{NLHS2}
\ee
\xxx{NLHS2}
\yyy{p. DHS-17, last box}


The exact state vector for $t \ge 0 $\/ is, from (\ref{Hilbertspace2}), (\ref{psi0HS2}) and the Schr\"{o}dinger equation,
\be
|\psi_{\cH_2}(t)\ra=c_1\exp\left(\frac{-iE_1 t}{\hbar}\right)|\cS_1\ra + c_1\exp\left(\frac{-iE_2 t}{\hbar}\right)|\cS_2\ra.\label{psitHS2}
\ee
\xxx{psitHS2}
\yyy{p. DHS-15, box 6}
So, the shortest time scale characterizing changes in $|\psi_{\cH_2}(t)\ra$\/ is
\be
\Delta t_{\cH_2}=2\pi \hbar/\max_i(|E_i|),\label{DeltatHS2}
\ee
\xxx{DeltatHS2}
\yyy{p. DHS-18, boxes 2-6}
in that $|\psi(t)\ra_{\cH_2}$\/ will be approximately constant over time intervals much smaller than (\ref{DeltatHS2}).

Using (\ref{psitodeltapsi}), (\ref{psi0HS2}) and (\ref{psitHS2}),
\be
\la\Delta \psi_{\cH_2}(\Delta t)|\Delta \psi_{\cH_2}(\Delta t)\ra^\frac{1}{2}=
          \left[2-2|c_1|^2\cos\left(\frac{E_1\Delta t}{\hbar}\right)-2|c_2|^2\cos\left(\frac{E_2\Delta t}{\hbar}\right)\right]^\frac{1}{2}.
\ee
If 
\be 
\Delta t \ll \Delta t_{\cH_2} \label{approxvalidHS2}
\ee
\xxx{approxvalidHS2}
\yyy{p. DHS-18, boxes 2-6}
then
\be
\la\Delta \psi_{\cH_2}(\Delta t)|\Delta \psi_{\cH_2}(\Delta t)\ra^\frac{1}{2}=
\left[|c_1|^2E_1^2 + |c_2|^2E_2^2\right]^\frac{1}{2}\frac{\Delta t}{\hbar},
\ee
so the norm of the change in the state vector during the minimum possible time interval $\Delta t_{min}$\/ (assumed, as discussed above, to satisfy
(\ref{approxvalidHS2})) is greater than or equal to the limit (\ref{NLHS2}) deduced from the RQSL:
\be
\la\Delta \psi_{\cH_2}(\Delta t_{min})|\Delta \psi_{\cH_2}(\Delta t_{min})\ra^\frac{1}{2} \ge {\rm NormLim}_{{\rm RQSL} - \cH_2}
\ee

\subsection{Ideal measurement} \label{SubsectionIdealMeasurement}

We consider the simplest ideal measurement\ 
model, perhaps more appropriately described as  a model of a ``detector'' rather than a ``measuring device.''  An observer (a human or an apparatus) starts in a ready state $|\cO_1\ra$\/ and measures a two-state system
that starts and remains in its initial state.
If the  observer finds that the measured system is in state 
$|\cS_1\ra$\/ the observer  remains in the ready state; if she finds the measured system to be in state  $|\cS_2\ra$\/ the observer transitions to a state 
$|\cO_2\ra$\/.

In detail: The Hilbert space $\cH_D$\/ is spanned by the four vectors $|\cS_i\ra \otimes |\cO_j\ra,$\/ $i,j=1,2$\/, where
\be 
\la \cS_i|\cS_j\ra=\la \cO_i|\cO_j\ra=\delta_{i,j}, \hspace*{1cm} i,j=1,2. \label{HDstates}
\ee
\xxx{HDstates}
\yyy{p. DHS-20, first line}
The Hamiltonian is
\be
\wh{H}_{\cH_D}=\wh{P}^{\cS_2} \otimes \wh{h}^\cO \label{HamiltonianHD}
\ee
\xxx{HamiltonianHD}
\yyy{p. DHS-20, box 2}
where 
\be
\wh{P}^{\cS_2}=|\cS_2\ra\la \cS_2| \label{projectorS2}
\ee
\xxx{projectorS2}
\yyy{p. DHS-20, box 3}
and
\be
\wh{h}^\cO=i\kappa\left(|\cO_2\ra\la \cO_1| - |\cO_1\ra\la \cO_2|\right) \label{hO}
\ee
\xxx{hO}
\yyy{p. DHS-20, box 4}
The state vector describing the observer and measured system at the initial time $t=0$\/ has the observer in the ready state and uncorrelated with the measured system:
\be
|\psi_{\cH_D}(0)\ra=\left(c_1|\cS_1\ra + c_2|\cS_2\ra\right)\otimes|\cO_1\ra,\hspace*{1cm}|c_1|^2+|c_2|^2=1. \label{psiHD0}
\ee
\xxx{psiHD0}
\yyy{p. DHS-20, box 1}
Using (\ref{DeltaHdef}), (\ref{NLRQSLdef}) and (\ref{HDstates})-(\ref{psiHD0}), we find the 
lower limit on the norm of the change in the state vector from $t=0$\/ to $t=\Delta t_{min}$\/ to be
\be
{\rm NormLim}_{{\rm RQSL} - \cH_D} = \frac{|c_2|\kappa \Delta t_{min}}{\hbar}. \label{NLHSD}
\ee
\xxx{NLHSD}
\yyy{p. DHS-22, boxes 1,3}

The exact state vector for $t \ge 0$\/ is, from (\ref{HamiltonianHD})-(\ref{psiHD0}) and the Schr\"{o}dinger equation,\footnote{Note from (\ref{psitHD}) that at time $t_{meas}=\pi \hbar/(2\kappa)$\/ the ideal measurement has completed, in that the observer and observed-system states have become perfectly correlated.  This confirms that the choice of Hamiltonian (\ref{HamiltonianHD}) and initial state (\ref{psiHD0}) indeed models an ideal measurement. Of course $t_{meas}$\/ is much longer than $\Delta t_{min}$\/; 
see (\ref{DeltatHD}), (\ref{approxvalidHD}).}
\be
|\psi_{\cH_D}(t)\ra=c_1|\cS_1\ra|\cO_1\ra+c_2|\cS_2\ra\left[\cos\left(\frac{\kappa t}{\hbar}\right)|\cO_1\ra+\sin\left(\frac{\kappa t}{\hbar}\right)|\cO_2\ra\right] \label{psitHD}
\ee
\xxx{psitHD}
\yyy{p. DHS-26, box 1}
From (\ref{psitHD}) we see that the time scale characterizing changes in $|\psi_{\cH_D}(t)\ra$\/ is 
\be
\Delta t_{\cH_D}=\frac{2\pi \hbar}{|\kappa|}. \label{DeltatHD}
\ee
\xxx{DeltatHD}
We can arrive at the same conclusion by noting that the distinct eigenvalues of $\wh{H}_{\cH_D}$\/ are $0, \kappa$\/  and $-\kappa$\/, so
\be
\Delta t_{\cH_D}=2\pi \hbar/\max(|0|,|\kappa|,|-\kappa|). \label{DeltatHD2}
\ee
\xxx{DeltatHD2}
\yyy{p. DHS-30, last 4 boxes}
From (\ref{psitodeltapsi}), (\ref{HDstates}), (\ref{psiHD0}) and (\ref{psitHD}),
\be
\la\Delta \psi_{\cH_D}(t)|\Delta \psi_{\cH_D}(t)\ra^\frac{1}{2}=|c_2|\left[2-2\cos\left(\frac{\kappa t}{\hbar}\right)\right]^\frac{1}{2}.\label{exactnormHD}
\ee
\xxx{exactnormHD}
\yyy{p. DHS-27, box 4}
For $\Delta t$\/ satisfying
\be
\Delta t \ll \Delta t_{\cH_D} \label{approxvalidHD}
\ee
\xxx{approxvalidHD}
the norm of the change in the state vector, (\ref{exactnormHD}), becomes
\be
\la\Delta \psi_{\cH_D}(\Delta t)|\Delta \psi_{\cH_D}(\Delta t)\ra^\frac{1}{2}=\frac{|c_2|\kappa \Delta t}{\hbar}. \label{DeltatnormHD}
\ee
\xxx{DeltatnormHD}
\yyy{p. DHS-27, last box}
Taking $\Delta t$\/ to be the minimum possible time increment $\Delta t_{min}$\/, we see from (\ref{NLHSD}) and (\ref{DeltatnormHD})  that in this case the norm of the change in the state vector is equal to the limit coming from the RQSL:
\be 
\la\Delta \psi_{\cH_D}(\Delta t_{min})|\Delta \psi_{\cH_D}(\Delta t_{min})\ra^\frac{1}{2}= {\rm NormLim}_{{\rm RQSL} - \cH_D}.\label{NLRQSLHDcomparison}
\ee
\xxx{NLRQSLHDcomparison}

%
%
%
%
%
%

\section{Discussion}



While the arguments and results in \cite{BHZ05,CalmetHsu21,BHZ06,Hsu12,Hsu17,Hsu21,Rubin21} and the present paper suggest 
that the usual continuous Hilbert space is not the the proper arena in which to do quantum physics they do not  make this case conclusively, 
let alone 
point to 
a mathematical structure to serve as a replacement. 
Nevertheless Hilbert-space discreteness is an idea that 
deserves continued investigation, 
particularly for its potential to
explain one of the most recondite conceptual problems in physics and philosophy of science, that of the meaning and nature of probability.  

This issue is particularly salient for probability in the Everett interpretation of quantum mechanics.  In that theory, all outcomes of quantum processes occur in parallel, and since the inception of the theory physicists have struggled to understand the sense in which probabilities can be associated with outcomes 
if everything that can happen does happen \cite{Everett57,DeWitt72,Graham73,Okhuwa93,Kent90,Deutsch99,Barnumetal00,Rae09,Saunders10,Papineau10,Wallace10,GreavesMyrvold10,Kent10,Albert10,Price10,Zurek10,Schack10,Vaidman11,SebensCarroll18}.  

For example, when the ideal measurement of Sec.\ref{SubsectionIdealMeasurement} is complete, at time
 $t=t_{meas}=\pi \hbar/(2\kappa)$\/,
the state vector  (\ref{psitHD}) has the form 
\be
|\psi_{\cH_{D}}(t_{meas})\ra=c_1|\cS_1\ra|\cO_1\ra+c_2|\cS_2\ra |\cO_2\ra. \label{psitHDmeas}
\ee
\xxx{psitHDmeas}
In the standard (single-outcome, non-Everett) 
interpretation of quantum mechanics commonly presented in textbooks, 
a post-measurement state such as (\ref{psitHDmeas}) 
is replaced stochastically by 
one of the states corresponding to the possible outcomes of the measurement, either
$|\cS_1\ra|\cO_1\ra$\/ 
(``system is in state 1, observer is in the ready state 1'') or  
$|\cS_2\ra |\cO_2\ra$\/
 (``system is in state 2, observer is in state 2''), 
with the respective probabilities for each of these outcomes occurring being the absolute-squared amplitudes $|c_1|^2$\/, $|c_2|^2$\/.  This is termed ``reduction of the state vector''\cite[p. 236]{Ballentine98}.

In the Everett interpretation, on the other hand, there is no reduction.  The two terms on the right-hand side  of (\ref{psitHDmeas}) represents two Everett worlds. In one of these worlds the  outcome $|\cS_1\ra|\cO_1\ra$\/  has occurred, in the other the outcome $|\cS_2\ra |\cO_2\ra$\/ has occurred.

What can it possibly mean to say that one outcome is more or less probable than the other when in fact {\em both}\/ have occurred?\footnote{It is precisely the existence of multiple outcomes that is critical in enabling  the Everett interpretation to evade Bell's theorem and the resulting conclusion that quantum mechanics is nonlocal; see \cite[Sec. 1.1]{Rubin11} and references therein.}  This question is referred to as the ``incoherence problem''\cite{Wallace12}. 

The incoherence problem presents itself in a particularly acute form if we consider a situation in which one or more of the 
 outcomes is a  world corresponding to a state of affairs that is assigned such small probability by the standard quantum interpretation that we would feel justified in neglecting even the possibility of its occurring. For example, such a world might be one in which large violations of the second law of thermodynamics occur, with  heat flowing from colder to warmer bodies, gases and liquids spontaneously unmixing, etc.  Such worlds are termed ``maverick worlds''\cite{DeWitt70}.   Of course our intuition about the nature of probability leads us to feel that we understand why  any outcome with exceedingly small probability can be neglected.\footnote{Kolmogorov: ``We apply the theory of probability to the actual world of experiments in the following manner\ldots If P($A$\/) is very small, one can be practically certain that when conditions [for a repeatable experiment] are realized only once, the event $A$\/ would not occur at all\cite[pp. 3-4]{Kolmogorov56}.''} But, in the absence of a meaning we can assign to probability small or large, we are unable to explain why maverick outcomes are not features of our day-to-day experience. 

Incorporating Hilbert-space discreteness into Everett quantum mechanics provides  
a resolution to the problem of maverick worlds.  It implies that there is a minimum norm to state vectors, and one can then argue that maverick worlds, having norms below this minimum, are simply absent from any superposition of  outcomes\cite{BHZ06}. 

This might seem to still leave open the issue of how to assign meaning to probabilities {\em larger} than, say, those involving violation of the second law of thermodynamics---i.e., the probabilities encountered in the vast majority of physical situations.  But in fact, preclusion of state-vector components with norms below a certain minimum holds the potential of giving meaning to 
these probabilities as well.
%
%
%
For those probabilistic phenomena that are not precluded but that in fact may or may not be observed to occur, probability may be identified with  
subjective experience of a particular kind.  In the words of  de Finetti, one of the most prominent proponents of this subjective interpretation of  probability: ``What do we mean when we say, in ordinary language, that an event is more or less probable? We mean that we would be more or less surprised to learn that it has not happened\cite[p. 174]{deFinetti89}.''   Furthermore, 
there is a substantial body of {\em experimental}\/ evidence (see \cite[Sec. 6.4]{Rubin21} and references therein)  that the subjective experience of probability  is at its base a result of biological evolution\cite{Koopmans40}, \cite[Sec. 6.1]{Rubin21}. And it can be argued  that the  biological evolution, over generation upon generation, of subjective-probabilistic expectations fit to maverick worlds will be of sufficiently small norm to be precluded, leaving only organisms with subjective expectations matching our familiar notions of probability and statistics.
In a nutshell, ``preclusion explains evolution which then explains subjective probability\cite[p. 21]{Rubin21}.''
A proof-of-concept quantum-mechanical model of this process is presented in \cite[Secs. 6.2, 6.3]{Rubin21}. 

In addition to thus accounting for our experiences of probabilistic phenomena within the completely-deterministic Everettian framework, Everett quantum mechanics with preclusion of states below a minimum norm  
addresses  longstanding interpretational problems of probability {\em per se}\/\cite{SEPprobability}.  Along with purely-subjective theories of probability, this theory has the virtue of leaving no ambiguity as to what is meant by probability; namely, the subjective reactions of organisms to probabilistic  situations.   At the same time, the objective feature of preclusion 
allows it to provide 
{\em explanations}\/ 
for physical phenomena in general and for biological evolution in particular, something which purely subjective theories are incapable of doing.  An explanation should at the very least be 
``a statement of what is there in reality, and how it behaves and how that accounts for the explicanda\cite{Deutsch16}.'' So, the  explanation of  objectively-existing phenomena (including subjective probability\footnote{Subjective reactions are themselves objectively-existing phenomena.  ``The subjective opinion, as something known by the individual under consideration is, at least in this sense, something objective and can be a reasonable subject  of a rigorous study''\cite[p. 5]{deFinetti17}.}) must itself be something that exists objectively (in this theory, preclusion and the norms of quantum states).

For a detailed exposition of probability in the version of Everett quantum theory with preclusion outlined here, see \cite[Secs. 6.1, 7, 8]{Rubin21}. Other approaches to solving the Everett probability problem via discrete Hilbert space are presented in 
\cite{BHZ06,Hsu12, Hsu17, Hsu21}.  
Much more work needs to be done to place minimum Hilbert space norms on a concrete foundation, but the completion of such a program should prove highly consequential.



\section*{Acknowledgments}

I would like to thank Jianbin Mao, Jacob A. Rubin and Allen J. Tino for helpful discussions.

\end{document}